OpenStudy (anonymous):

Given the function f(x) = 0.3(4)^x, what is the value of f^−1 (6)? 0.360 1.242 2.161 2.312

OpenStudy (lucaz):

@mathmale ?

OpenStudy (mathmale):

Goal: Find the inverse function of f(x) = 0.3(4)^x f(x) = 0.3(4)^x Method: 1) replace that f(x) with y: y = 0.3(4)^x 2) interchange x and y 3) solve the resulting equation for y 4) replace the y in your final answer with \[f ^{-1}(x).\]

OpenStudy (lucaz):

 is that correct?